Suppose f(x) = 1. 5 đť‘Ą for -1 < x <1. Determine the mean and variance of X


1
Expert's answer
2022-06-14T12:05:39-0400

Check


"\displaystyle\int_{-\infin}^{\infin}f(x)dx=\displaystyle\int_{-1}^{1}1.5x^2dx"

"=[0.5x^3]\begin{matrix}
 1 \\
-1
\end{matrix}=1, True"


"\mu=E(X)=\displaystyle\int_{-\infin}^{\infin}xf(x)dx=\displaystyle\int_{-1}^{1}1.5x^3dx"

"=[\dfrac{1.5}{4}x^4]\begin{matrix}
 1 \\
-1
\end{matrix}=0"

"E(X^2)=\displaystyle\int_{-\infin}^{\infin}x^2f(x)dx=\displaystyle\int_{-1}^{1}1.5x^4dx"

"=[0.3x^5]\begin{matrix}
 1 \\
-1
\end{matrix}=0.6"

"Var(X)=E(X^2)-(E(X))^2"

"=0.6-(0)^2=0.6"